OpenStudy (anonymous):

Find the equation of the line that contains the point (3, -5) and is perpendicular to the line that contains (13, 3) and (-1, -3) -this one has stumped me for a while, help?

OpenStudy (anonymous):

perpendicular means slope is negative reciprocal, so find the slope of the line between (12,3) and (-1,-3) i get \[\frac{3+3}{13+1}=\frac{6}{14}=\frac{3}{7}\] thus perpendicular line will have slope \[-\frac{7}{3}\]

OpenStudy (anonymous):

now use almighty point - slope formula to find the equation of the line \[y+5=-\frac{7}{3}(x-3)\] \[y+5=-\frac{7}{3}x+7\] \[y=-\frac{7}{3}x+2\]